OpenStudy (anonymous):

Anyone familiar with Homogeneous 3D Coordinates?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Probably--homogenous means many things in mathematics. Post the question let's see. What class are you in? I have had linear algebra and vector calculus; I'm taking multivariate calc and diffEQ now.

OpenStudy (anonymous):

I am in linear algebra

OpenStudy (anonymous):

how do you like multivariable calc? That was a fun class

OpenStudy (anonymous):

Did you guys talk about applications to computer graphics in linear algebra?

OpenStudy (anonymous):

Fun so far--it's Calc III, and we've just finished our vector calculus test. Partial derivatives and multiple integrals are easy--they've come up in physics and diffEQ. I've heard it's difficult to visualize, so we'll see how hard it is. I found linear algebra to be extremely fun--but the notation and vocabulary isn't standard, which is why lots of context is needed.

OpenStudy (anonymous):

We didn't talk about it explicitly, but we talked about converting co-ordinates in N dimensions. I had planned to do a presentation on the applications of computer graphics, but I couldn't find any resources I liked, so I did graph theory instead.

OpenStudy (anonymous):

Haha I am opposite, vector calc was easy for me but linear is another story.

OpenStudy (anonymous):

I am a visual learner and i like applied mathematics instead of dealing with numbers, theory, proofs, etc/

OpenStudy (anonymous):

I am in the processing of writing a report on computer graphics. I am just having a bit of hard time understand Homogeneous 3D Coordinates

OpenStudy (anonymous):

That's one of the reason's studying N dimensions is easier than studying 3 dimensions. You have to visualize things on a very weird, limited level. 2 dimensions with parameters, essentially. I will look through my book and see what I can find.

OpenStudy (anonymous):

This isn't my paper I wrote but he explains things the same as my textbook does

OpenStudy (anonymous):

when he uses X=x/h, Y=y/h, and Z=z/H why is he dividing everything by half? Why are we scaling all the coordinates by half?

OpenStudy (anonymous):

No particular reason, it's just an example. That's something that's very difficult to get used to in advanced mathematics--someone will say "Let x=sqrt(5000)/pi^9" or whatever. The key is to just keep going and realize that in math, I'm allowed to define anything I want to. I haven't read any further, but if I remember the computer graphics portion I read, we throw 3 dimensional objects into 4 dimensions so that we can pull out the parts of the object we want in R^3. The book hasn't yet said anything about homogenous co-ordinates. They will probably be mentioned again later in the book--or paper.

OpenStudy (anonymous):

you familiar with perspective projection?

OpenStudy (anonymous):

Haven't heard of it, but It sounds like vector projection in r^3. Just transforming all the points on a surface closest to a plane onto that plane? My class was very focused on the theoretical aspects. Vector spaces, inner product spaces--lots of proofs. I am reading later into the paper atm.